OpenStudy (anonymous):

Seamus invested an average of $325 per month since age 40 in various securities for his retirement savings. His investments averaged a 5.5% annual rate of return until he retired at age 68. Given the same monthly investment and rate of return, how much more would Seamus have in his retirement savings had he started investing at age 30? Assume monthly compounding. $123,435.00 $214,500.00 $240,962.89 $268,913.99

OpenStudy (phi):

\[ \text{ future value }= \text{pymt } \cdot \left( \frac{ (1+i)^n - 1}{i} \right) \] where i is the monthly interest rate, n is the total number of months, and pymt is the payment= $325 per month you can calculate the future value for 18 years (n= 28*12 months) and then for 28 years (n= 38*12 months)